4. Simplify each expression. All variables represent positive real numbers 3/24x15y4 a. 3)(x + 6)3 b. 3/7
Answer #1

Note: of (c +673 2+1 96 = sta at6 Ans 5 3 24 x 5 yt 2124 ho 2 12 2 1/3 2 2²x3=24 (23x3)<<!]44 (313 Note: 2x3s persona y 92 x